1a Determine the possible number of macroscopic states for 15 molecules in two intersecting cells of equal volume.

b.detrmine the possible number of microscopic states.

c,determine the possible number of microscopic states for 8 molecules in cell A and 7 molecules in cell B.




1
Expert's answer
2021-07-20T15:18:10-0400

1.

"g_i=15,n_i=2"

Macroscopic state

gCn

Marcroscopic state

"\frac{g_i!}{g_i-n_i ! n_i}"

Put value

macroscopic state

"\frac{15!}{(15-2)! 2!}=105"

2.

Macroscopic format

"n_iC^{g_i}"

"=\frac{g_i!}{g_i-n_i ! n_i}"

3

"g_i=7,g_i=8\\n_i=1,n_i=1"

"C_1^7\times C_1^8"

"\frac{7!}{6!\times 1!}\times\frac{8!}{7!\times 1!}=7\times8=56"